OpenStudy (anonymous):

MEDALS For mega help -Is the following statement true or false? If it is true, explain your reasoning.

OpenStudy (muzzack):

false

OpenStudy (anonymous):

yes like @Muzzack said the answer false because if u substitute some values for a and b such as 3 and 4 respectively u get 2(3)*4=3(2)+4(2) which is a false statement. In order for the statement to be true the left hand side has to = the right hand side .. the left side is saying 6*4 which is 24 and the right side is saying 6+8 which is 14 so thats not =... to make it equal the right side should say 3(2)+4(4.5) then u would have 24=24